-10 points DevoreSts:8 14 E.008 An article focuses on the existence of any relationship between the date of patient admission for treatment of elcoholism and the patients birthday. Assuming 365-dey year [i.e. excluding leap year), in the absence of any relation, a petients admission date is equally likely to be any one of the 365 possible days. The investigators established four different admission categories: (1) within 7 days of birthday, [2) betreen 8 and 30 days, inclusive, from the birthday, (3) between 31 and 90 days, inclusive from the birthday, and (4) more than 90 days from the birthday. A sample of 196 patients gave observed frequencies of 12, 22, 65, and 97 for catagories 1, 2, 3, and 4, respectivaly. State and test the relevant hypothezes using a significanco laval of0.01. My Notes Ask Your 46 365 at least one P-Po 184 46 120 184 ,: at least one P-Po State the rejection regian and compute the test statistic value. Round your answers to three decimal places.) rajection ragion test statistic
Answer #1

for hypothesis: last option is correct

here degree of freedom =categories-1=4-1=3

for 3 degree of freedom ; rejection region \chi ^{2} >=11.345

applying chi square goodness of fit:\

observed Expected Chi square
category Probability(p) Oi Ei=total*p R2i=(Oi-Ei)2/Ei
Monday    3/73 12.000 8.055 1.932
Tuesday 46/365 22.000 24.701 0.295
Wednesday 24/73 65.000 64.438 0.005
Thursday 184/365 97.000 98.805 0.033
total 1.000 196 196 2.266

test statistic \chi ^{2} =2.266